OpenStudy (anonymous):

Suppose you earn $7.25 per hour working part-time for a florist. What is the minimum number of hours you must work to earn at least $125.00? Write an inequality to represent this situation and solve..

OpenStudy (mayankdevnani):

7.25 for working 1 hour 1 for working 1/7.25 hours 125 for working (1/7.25)*125 = 125/7.25 = 17.24 hours so to earn at least 125 she has to work 18 hours @bstarr

OpenStudy (anonymous):

thanks.